# A Data-driven Typology of Vision Models from Integrated Representational Metrics

### v2

@@ 2.3 Separation Ability Metrics
The contrastive ratio quantifies the relative separation between intra-family and inter-family similarities. We consider the similarity values between different models within the same family and take the average of them to obtain $\mu_{\text{within}}$, and the similarities between models from two model families and obtain the average inter-family $\mu_{\text{between}}$. The ratio is then defined as $CR=(\mu_{\text{within}}-\mu_{\text{between}})/(\mu_{\text{within}}+\mu_{\text{between}})$. A value approaching 1 suggests strong within-family coherence relative to cross-family similarity; a value approaching 0 suggests no difference, and a negative value implies that inter-family similarity exceeds intra-family similarity.
D-Prime (Bo et al., 2024).
@@
Similarly to contrastive ratio but considering the variance, the D-Prime ($d^{\prime}$) also quantifies the separation between intra-family and inter-family similarity distributions. It is defined as $d^{\prime}={(\mu_{\text{within}}-\mu_{\text{between}})}/{\sqrt{0.5(\sigma^{2}_{\text{within}}+\sigma^{2}_{\text{between}})}}$, where $\mu$ and $\sigma^{2}$ denote the mean and variance of the respective distributions. Higher values indicate tighter clustering within a family and greater spread across families, reflecting stronger separability.
Silhouette Score (Rousseeuw, 1987).
@@
For each model $i$, we compute the average distance $a(i)$ to all other models in the same family and the average distance $b(i)$ to models in the other family. The silhouette value is then $s(i)=({b(i)-a(i)})/{\max\{a(i),b(i)\}},s(i)\in[-1,1]$. Values near $1$ indicate that the model is well grouped with its own family, values near 0 suggest boundary placement, and negative values imply greater similarity to another family. The overall silhouette score is obtained by averaging $s(i)$ across all models.
@@ 2.4 Similarity Network Fusion
Next, we sought to reconcile the results across the different evaluation metrics. As elaborated in Section 2.3, each metric captures distinct aspects of model representations and varies in its ability to differentiate between model families. To integrate these metrics, we adopt a unified approach inspired by Similarity Network Fusion (Wang et al., 2014; Markello, 2020). Let $n$ be the number of models, and $\mathbb{V}$} be the set of the representational metrics. For each metric $v\in{\mathbb{V}}$, we can get a similarity matrix $\mathbf{S}^{v}\in\mathbb{R}^{n\times n}$, where each entry $\mathbf{S}^{v}_{ij}$ measures the similarity between the model $i$’s representation and model $j$’s according to metric $v$, as described in Section 2.2. Then, for each metric $v$, we first convert pairwise scores into a dissimilarity matrix $\mathbf{Q}^{v}$ as this equation, $Q^{v}_{ij}=1_{\mathrm{i\neq j}}\!\left(1-(S^{v}_{ij}+S^{v}_{ji})/{2}\right)$. We then build an affinity $\mathbf{W}^{v}\in\mathbb{R}^{M\times M}$ with a scaled exponential kernel:
$\mathbf{W}^{v}(i,j)=\frac{1}{\sqrt{2\pi(\sigma^{v}_{ij})^{2}}}\exp\left(-\frac{(\mathbf{Q}^{v}_{ij})^{2}}{2(\sigma^{v}_{ij})^{2}}\right),$ with  $\sigma^{v}_{ij}=\mu\cdot\frac{\overline{\mathbf{Q}}^{v}(i,N_{i})+\overline{\mathbf{Q}}^{v}(j,N_{j})+\mathbf{Q}^{v}_{ij}}{3}$.
@@
$\mathbf{C}^{v}_{ii}=\sum_{j}\mathbf{W}^{v}_{ij}$,  $\widetilde{\mathbf{W}}^{\,v}=(\mathbf{C}^{v})^{-1}\mathbf{W}^{v}$,  $\widehat{\mathbf{W}}^{\,v}=\tfrac{1}{2}(\widetilde{\mathbf{W}}^{\,v}+(\widetilde{\mathbf{W}}^{\,v})^{\top})$,
@@
$\mathbf{P}^{(v)}_{t+1}=\mathcal{B}_{\alpha}\!\left(\mathbf{S}^{(v)}\left(\tfrac{1}{|{\mathbb{V}}|-1}\sum_{u\neq v}\mathbf{P}^{(u)}_{t}\right)\mathbf{S}^{(v)\top}\right)$,  $\mathcal{B}_{\alpha}(\mathbf{X})=\tfrac{1}{2}(\mathbf{X}+\mathbf{X}^{\top})+\alpha\mathbf{I}$
@@
$\mathbf{P}=\frac{1}{|{\mathbb{V}}|}\sum_{v\in{\mathbb{V}}}\mathbf{P}^{(v)}_{T}$,  $\mathbf{D}_{ii}=\sum_{j}\mathbf{P}_{ij}$,  $\widetilde{\mathbf{P}}=\mathbf{D}^{-1}\mathbf{P}$,  $\widehat{\mathbf{P}}=\tfrac{1}{2}\big(\widetilde{\mathbf{P}}+\widetilde{\mathbf{P}}^{\top}+\mathbf{I}\big)$.
@@ 3 Results
3.1 Divergent versus convergent dimensions across model families
To understand which aspects of learned representations are universally shared across vision models and which constitute family-specific signatures, we systematically evaluate multiple representational facets. Each metric captures a distinct dimension of representational structure, allowing us to identify where model families converge versus diverge. To assess family discriminability, we quantified the difference between within-family and across-family representational similarities using the separability measures introduced above. The results presented here are based on ImageNet; consistent patterns were also observed on other datasets (Appendix C). Our systematic evaluation reveals that different representational dimensions show markedly different patterns of convergence versus divergence across model families (Figures 23,  D.1,  D.2,  D.3). This variation suggests that while some computational strategies are universally adopted across architectures and training paradigms, others constitute distinctive signatures of specific model families.
@@ 3.1 Divergent versus convergent dimensions across model families
Metrics that preserve representational geometry or unit-level tuning properties demonstrate the strongest ability to discriminate between model families. RSA, which captures representational geometry, achieves the highest separability with a $d^{\prime}$ of 3.95 and silhouette coefficient of 0.56—indicating that geometric organization or relational structure—how models arrange points in representation space—constitutes a strong family-specific signature. Linear CKA also shows strong discrimination ($d^{\prime}$ = 3.91), which aligns with recent theoretical work showing that centered RSA and linear CKA are mathematically equivalent when appropriate centering is applied (Williams, 2024). Similarly, SoftMatch, which preserves individual unit tuning while mapping two representations, shows robust discrimination ($d^{\prime}$ = 3.59, silhouette = 0.30). Even supervised and self-supervised variants within the same architecture family (particularly CNNs) are reliably separated by these metrics, demonstrating that the training paradigm fundamentally shapes the geometry and tunings of individual neurons, such that metrics diagnostic of these representational facets achieve good separation. These properties thus constitute the unique representational “fingerprints” of model families.
@@
Interestingly, Procrustes alignment—which allows orthogonal transformations—shows intermediate discrimination ($d^{\prime}$ = 2.96), falling between SoftMatch and Linear Predictivity. This reveals a clear pattern among mapping-based metrics: discriminability decreases monotonically as the transformations become more flexible (SoftMatch >> Procrustes >> Linear Predictivity). The constraints imposed by less flexible mappings appear to preserve family-specific signatures that are lost when arbitrary linear transformations are allowed.
@@
Figure 2: Model-family separability on ImageNet under $d^{\prime}$, silhouette score and contrastive ratio. Columns correspond to nine similarity metrics, including two fusion-based methods (SNF, average) and seven commonly used representational metrics (Distinct aspects of representation emphasized by each metric are shown in the bracketed text). Fusion-based metrics consistently yield higher scores, highlighting their effectiveness in capturing family-level distinctions.
@@
In contrast, metrics capturing linearly accessible information show substantially weaker discrimination between families. Linear Predictivity demonstrates the lowest separability among direct mapping-based metrics ($d^{\prime}$ = 2.09, silhouette = 0.14), while CCA-based metrics (PWCCA: $d^{\prime}$ = 1.55; SVCCA: $d^{\prime}$ = 1.02) show even weaker family separation. The weak discrimination of CCA-based metrics is particularly revealing. CCA identifies maximally correlated linear projections between representations, finding shared subspaces that are invariant to invertible linear transformations. CCA loads on the linear-accessibility facet: it detects shared linearly decodable subspaces but, unlike RSA/CKA or Procrustes/SoftMatch, it does not constrain or preserve representational geometry or tuning. The invariance of CCA to linear transformations, which makes it powerful for finding shared structure across superficially dissimilar representations, also makes it insensitive to the geometric and topological features that distinguish model families.
@@
The theoretical relationships among these metrics help explain the discrimination hierarchy. RSA and Linear CKA are mathematically equivalent under appropriate centering (Williams, 2024) and both preserve the geometric structure of representations—they compare how similarly models organize their representation spaces without fitting any transformation. In contrast, the mapping-based metrics show decreasing discrimination as they allow increasingly flexible transformations: SoftMatch permits only permutations that preserve individual unit correspondences, Procrustes allows orthogonal transformations (rotations and reflections), while CCA searches for optimal linear projections that maximize correlation. Linear Predictivity provides the most flexibility, allowing any linear transformation that minimizes prediction error.
@@
The weak discrimination of metrics that assess linearly accessible information might suggest that this aspect of representation is more consistent across model families than geometric organization. Whether this reflects convergent computational strategies, methodological limitations of these metrics, or task-imposed constraints remains an open question.
3.2 Integration achieves superior model family discrimination
Figure 3: Mean model-family separability on ImageNet, evaluated using $d^{\prime}$, silhouette score, and contrastive ratio. Fusion-based metrics (SNF, Average) outperform individual similarity metrics across all datasets, with SNF yielding the most consistent and robust separation. Scores are shown in their native scales and are not directly comparable across measures.
@@ 3.2 Integration achieves superior model family discrimination
Critically, our SNF approach, which integrates information across all representational dimensions, achieves dramatically superior family separation compared to any single metric. SNF attains a $d^{\prime}$ of 11.84—nearly three times higher than the best-performing single measure—and consistently outperforms all baselines across separation criteria. Importantly, as shown in Figure 2, SNF maintains high and balanced discrimination across nearly all family pairs. By contrast, individual metrics often exhibit uneven performance, separating some families while failing or for others.
@@
To test whether SNF recovers the shared structure across metrics rather than replicating any single one, we try to quantify the intermetric agreement. Specifically, to ensure comparability across metrics, we symmetrize every similarity matrix by averaging it with its transpose, remove the diagonal (self-similarity), vectorize the remaining upper-triangle entries, and then compute correlation between vectors. As shown in Figure E.1, geometry-preserving metrics (RSA, SoftMatch, CKA) show strong mutual agreement, mapping-based metrics (Procrustes, Linear Predictivity) are moderately aligned, and CCA-variants highly correlate with each other but less agree with other metrics. SNF aligns only moderately with any single metric and is clearly distinct from simple averaging, indicating it fuses complementary facets instead of collapsing to one metric.
To further validate the SNF-derived similarity structure, we incorporate analyses based on biological neural data (Fig. I.1I.2). Using anatomically ordered visual areas across subjects, we find that SNF more accurately recovers expected relationships such as the alignment of homologous regions across subjects and the established ventral-stream hierarchy than any individual metric.
3.3 A data-driven typology of vision models
Having shown that integrated representational signatures provide the most comprehensive characterization of model families, we next use the SNF-fused similarity matrix to derive a data-driven typology. This typology reveals how models cluster according to their representational processing, moving beyond surface-level groupings defined by architecture or training paradigm.
@@ 3.3 A data-driven typology of vision models
As a baseline, we first examine clustering results from individual metrics (Figure 4, F.2, F.4, F.6). We observe distinct patterns across the metrics: measures such as Linear Predictivity and Procrustes, tend to produce relatively uniform similarity values across a wide range of models, resulting in diffuse, non-distinct clusters. SoftMatch, which emphasizes the geometric alignment of individual units, struggles to clearly separate models that are neither CNN-based nor supervised. PWCCA and SVCCA produce noisy similarity matrices in which, despite dendrogram reordering, no strong or coherent clustering structure is apparent. RSA reveals some clustering—most notably a separation between CNNs and Transformers—but the partitions appear quite diffuse. Even simple averaging of normalized metric values fails to produce sharply defined clusters. Together, these results highlight the limitations of single metrics: each emphasizes a different facet of representational similarity, leading to clustering patterns that are fragmented, noisy, or inconsistent across metrics.
Figure 4: Hierarchical clustering of models using three functionally distinct representational similarity metrics and a similarity-metric–averaging baseline (see Fig. F.1 for additional metrics). Clustering is performed with average linkage and optimal leaf ordering, based on induced distances (11- similarity score). Rows and columns are deliberately re-ordered to match the leaf ordering produced by the clustering algorithm. Lighter colors indicate higher similarity; diagonal entries (self-comparisons) are omitted.
Figure 5: SNF-based clustering reveals that models naturally group by architecture and supervision regime. Supervised CNNs and ViTs each form distinct clusters; hybrid models (ConvNeXt, Swin) and MAE ViTs cluster together; and self-supervised models (e.g., DINO ViTs, self-supervised CNNs) form coherent groups. The heatmap shows the SNF-fused similarity matrix reordered by leaf ordering. Leaf labels are colored by the cluster (formed by SNF) they belong to; dendrogram cuts yield up to six flat clusters aligned with canonical categories.
@@
In contrast, hierarchical clustering of the fused similarity matrix reveals well-defined groupings that both confirm expected relationships and uncover surprising organizational principles (Figure 5, F.3, F.5, F.7)). To quantitatively assess the fidelity of these clusters, we compute the cophenetic correlation coefficient (CCC) measuring how well the clusters preserves the original pairwise similarities (Figure G.1). Our SNF-based metric achieves one of the highest CCC of 0.982 among all metrics indicating that the clusters formed by the fusion matrix more faithfully reflect the underlying similarity structure among models. These results provide additional evidence that the SNF produces more meaningful and robust clusterings compared to individual baseline metrics, providing confidence in the typology.
Next, we examinine the clusters derived from the fusion matrix (Figure 5). Our typology reveals a complex interplay between architecture and training paradigm in determining representational structure. While some clusters align with architectural expectations—supervised ResNets group together, as do supervised VGGs and supervised ViTs—the most striking finding is that training paradigm can override architectural boundaries. All self-supervised models, regardless of architecture, form a unified cluster that transcends the CNN-Transformer divide. Self-supervised ResNets (MoCo, DINO, SwAV, Barlow Twins) group more closely with self-supervised ViTs than with their supervised architectural siblings. This suggests that the computational strategies induced by self-supervised learning—whether through contrastive learning, self-distillation, or redundancy reduction—create a shared representational signature that dominates over architectural differences.
@@
Perhaps most surprisingly, hybrid architectures (ConvNeXt and Swin) cluster with MAE models despite their different design philosophies. ConvNeXts modernize CNNs with Transformer-inspired components, Swins introduce CNN-like biases into Transformers, and MAE employs masked reconstruction—yet they converge on similar representational structures. This convergence suggests that architectural modernization and masked reconstruction, though approaching from different angles, arrive at similar computational solutions. These unexpected groupings demonstrate the power of our data-driven approach: we discover that emergent computational strategies—shaped by the interplay of architecture, training objective, and task demands—determine representational structure. The typology thus reveals the “species” of vision models, defined not just by their surface characteristics but by how they fundamentally process and organize visual information.

### v4

@@ 2.3 Separability Metrics
Contrastive Ratio. The contrastive ratio quantifies the relative separation between intra-family and inter-family similarities. We consider the similarity values between different models within the same family and take the average of them to obtain $\mu_{\text{within}}$, and the similarities between models from two model families and obtain the average inter-family $\mu_{\text{between}}$. The ratio is then defined as $CR=(\mu_{\text{within}}-\mu_{\text{between}})/(\mu_{\text{within}}+\mu_{\text{between}})$. A value approaching 1 suggests strong within-family coherence relative to cross-family similarity; a value approaching 0 suggests no difference, and a negative value implies that inter-family similarity exceeds intra-family similarity.
@@
D-Prime (Bo et al., 2024). Similarly to contrastive ratio but considering variances, the D-Prime ($d^{\prime}$) also quantifies the separation between intra-family and inter-family similarities. It is defined as ${(\mu_{\text{within}}-\mu_{\text{between}})}/{\sqrt{0.5(\sigma^{2}_{\text{within}}+\sigma^{2}_{\text{between}})}}$, where $\mu$ and $\sigma^{2}$ denote the mean and variance of the respective distributions. Higher values indicate tighter clustering within a family and greater spread across families.
@@
Silhouette Score (Rousseeuw, 1987). For each model $i$, we compute the average distance $a(i)$ to all other models in the same family and the average distance $b(i)$ to models in the other family. The silhouette value is then $s(i)=({b(i)-a(i)})/{\max\{a(i),b(i)\}},s(i)\in[-1,1]$. Values near $1$ indicate that the model is well grouped with its own family, values near 0 suggest boundary placement, and negative values imply greater similarity to another family. The overall silhouette score is obtained by averaging $s(i)$ across models.
@@ 2.4 Similarity Network Fusion
Next, we sought to reconcile the results across the different evaluation metrics. As elaborated in Section 2.3, each metric captures distinct aspects of model representations and varies in its ability to differentiate between model families. To integrate these metrics, we adopt a unified approach inspired by Similarity Network Fusion (Wang et al., 2014; Markello, 2020). Let $n$ be the number of models, and $\mathcal{V}$ be the set of the representational metrics. For each metric $v\in\mathcal{V}$, we can get a similarity matrix $\mathbf{S}^{v}\in\mathbb{R}^{n\times n}$, where each entry $\mathbf{S}^{v}_{ij}$ measures the similarity between the model $i$’s representation and model $j$’s according to metric $v$, as described in Section 2.2. Then, for each metric $v$, we first convert pairwise scores into a dissimilarity matrix $\mathbf{Q}^{v}$ as this equation, $Q^{v}_{ij}=1_{\mathrm{i\neq j}}\!\left(1-(S^{v}_{ij}+S^{v}_{ji})/{2}\right)$. We then build an affinity $\mathbf{W}^{v}\in\mathbb{R}^{M\times M}$ with a exponential kernel:
$\mathbf{W}^{v}(i,j)=\frac{1}{\sqrt{2\pi(\sigma^{v}_{ij})^{2}}}\exp\left(-\frac{(\mathbf{Q}^{v}_{ij})^{2}}{2(\sigma^{v}_{ij})^{2}}\right),$
$\sigma^{v}_{ij}=\mu\cdot\frac{\overline{\mathbf{Q}}^{v}(i,N_{i})+\overline{\mathbf{Q}}^{v}(j,N_{j})+\mathbf{Q}^{v}_{ij}}{3}$.
@@
$\mathbf{C}^{v}_{ii}=\sum_{j}\mathbf{W}^{v}_{ij}$, $\widetilde{\mathbf{W}}^{\,v}=(\mathbf{C}^{v})^{-1}\mathbf{W}^{v}$
$\widehat{\mathbf{W}}^{\,v}=\tfrac{1}{2}(\widetilde{\mathbf{W}}^{\,v}+(\widetilde{\mathbf{W}}^{\,v})^{\top})$,
@@
$\mathbf{P}^{(v)}_{t+1}=\mathcal{B}_{\alpha}\!\left(\mathbf{S}^{(v)}\left(\tfrac{1}{|\mathcal{V}|-1}\sum_{u\neq v}\mathbf{P}^{(u)}_{t}\right)\mathbf{S}^{(v)\top}\right)$
$\mathcal{B}_{\alpha}(\mathbf{X})=\tfrac{1}{2}(\mathbf{X}+\mathbf{X}^{\top})+\alpha\mathbf{I}$
@@
$\mathbf{P}=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\mathbf{P}^{(v)}_{T}$,  $\mathbf{D}_{ii}=\sum_{j}\mathbf{P}_{ij}$
$\widetilde{\mathbf{P}}=\mathbf{D}^{-1}\mathbf{P}$,  $\widehat{\mathbf{P}}=\tfrac{1}{2}\big(\widetilde{\mathbf{P}}+\widetilde{\mathbf{P}}^{\top}+\mathbf{I}\big)$.
@@ 3 Results
3.1 Which similarity metrics recover procedural differences between models?
To evaluate which dimensions of representational correspondence recover meaningful structure among models, we systematically compared a suite of representational similarity measures using the procedural criteria introduced above. Our goal is to assess how effectively each metric recovers known procedural structure, distinguishing models that differ in architecture or training paradigm while grouping procedurally matched models.
We quantify recovery of procedural structure using the separability measures described in Methods, which contrast within-family and across-family similarities. Results reported in this section are based on ImageNet representations; qualitatively similar patterns are observed across additional datasets (Appendix C). Across metrics, we observe substantial variation in the degree to which procedural structure is recovered (Figures 2C.1C.2C.3), indicating that different representational facets carry markedly different information about model families.
@@ 3.1 Which similarity metrics recover procedural differences between models?
Metrics that preserve representational geometry or unit-level tuning properties demonstrate the strongest ability to discriminate between model families. RSA, which captures representational geometry, achieves the highest separability with a $d^{\prime}$ of 3.95 and silhouette coefficient of 0.56, indicating that geometric organization or relational structure, i.e. how models arrange points in representation space, constitutes a strong family-specific signature. Linear CKA also shows strong discrimination ($d^{\prime}$ = 3.91), which aligns with recent theoretical work showing that centered RSA and linear CKA are mathematically equivalent when appropriate centering is applied (Williams, 2024). Similarly, SoftMatch, which preserves individual unit tuning while mapping two representations, shows robust discrimination ($d^{\prime}$ = 3.59, silhouette = 0.30). Even supervised and self-supervised variants within the same architecture family (particularly CNNs) are reliably separated by these metrics, demonstrating that the training paradigm fundamentally shapes the geometry and tunings of individual neurons, such that metrics diagnostic of these representational facets achieve good separation. These properties thus constitute the unique representational “fingerprints” of model families.
@@
Interestingly, Procrustes alignment, which allows orthogonal transformations, shows intermediate discrimination ($d^{\prime}$ = 2.96), falling between SoftMatch and Linear Predictivity. This reveals a clear pattern among mapping-based metrics: discriminability decreases monotonically as the transformations become more flexible (SoftMatch >> Procrustes >> Linear Predictivity). The constraints imposed by less flexible mappings appear to preserve family-specific signatures that are lost by arbitrary linear transformations.
@@
Figure 2: Model-family separability on ImageNet under $d^{\prime}$, silhouette score and contrastive ratio. Columns correspond to nine similarity metrics, including two fusion-based methods (SNF, average) and seven commonly used representational metrics (Distinct aspects of representation emphasized by each metric are shown in the bracketed text). The barplot on the right side of each row shows the mean model-family separability. Scores are shown in their native scales and are not directly comparable across measures.
Figure 3: Similar to Figure 2 but on the NSD dataset with human responses and excluding SVCCA and PWCCA.
@@
In contrast, metrics capturing linearly accessible information show substantially weaker discrimination between families. Linear Predictivity demonstrates the lowest separability among direct mapping-based metrics ($d^{\prime}$ = 2.09, silhouette = 0.14), while CCA-based metrics (PWCCA: $d^{\prime}$ = 1.55; SVCCA: $d^{\prime}$ = 1.02) show even weaker family separation. The weak discrimination of CCA-based metrics is particularly revealing. CCA identifies maximally correlated linear projections between representations, finding shared subspaces that are invariant to invertible linear transformations. CCA loads on the linear-accessibility facet: it detects shared linearly decodable subspaces but, unlike RSA/CKA or Procrustes/SoftMatch, it does not constrain or preserve representational geometry or tuning. The invariance of CCA to linear transformations, which makes it powerful for finding shared structure across superficially dissimilar representations, also makes it insensitive to the geometric and topological features that distinguish families.
@@
The theoretical relationships among these metrics help explain the discrimination hierarchy. RSA and Linear CKA are mathematically equivalent under appropriate centering (Williams, 2024) and both preserve the geometric structure of representations—they compare how similarly models organize their representation spaces without fitting any transformation. In contrast, the mapping-based metrics show decreasing discrimination as they allow increasingly flexible transformations: SoftMatch permits only permutations that preserve individual unit correspondences, Procrustes allows orthogonal transformations, while CCA searches for optimal linear projections that maximize correlation. Linear Predictivity provides the most flexibility, allowing any linear transformation that minimizes prediction error.
@@
The weak discrimination of metrics that assess linearly accessible information might suggest that this aspect of representation is more consistent across model families than geometric organization. Whether this reflects convergent computational strategies, methodological limitations of these metrics, or task-imposed constraints remains open.
3.2 Integration achieves superior discrimination
@@
Critically, our SNF approach, which integrates information across all representational dimensions, achieves dramatically superior family separation compared to any single metric. SNF attains a $d^{\prime}$ of 12.42—nearly three times higher than the best-performing single measure—and consistently outperforms all baselines across separation criteria. Importantly, as shown in Figure 2, SNF maintains high and balanced discrimination across nearly all family pairs. By contrast, individual metrics often exhibit uneven performance, separating some families while failing or for others.
@@
To test whether SNF recovers the shared structure across metrics rather than replicating any single one, we try to quantify the inter-metric agreement. Specifically, to ensure comparability across metrics, we symmetrize every similarity matrix by averaging it with its transpose, remove the diagonal (self-similarity), vectorize the remaining upper-triangle entries, and then compute correlation between vectors. As shown in Figure D.1, geometry-preserving metrics (RSA, SoftMatch, CKA) show strong mutual agreement, mapping-based metrics (Procrustes, Linear Predictivity) are moderately aligned, and CCA-variants highly correlate with each other but less agree with other metrics. SNF aligns only moderately with any single metric and is clearly distinct from simple averaging, indicating it fuses complementary facets instead of collapsing to one metric.
3.3 Which similarity metrics recover known anatomical–functional structure in visual cortex?
We next evaluate representational similarity metrics using neural data from the Natural Scenes Dataset (NSD), asking whether the patterns observed in artificial models generalize to biological visual systems. We compare representational correspondences across cortical regions using responses to a shared set of natural images, focusing on cross-subject comparisons. Comparisons across different regions within the same subject are excluded due to the spatial blurring inherent to fMRI data, which can artificially inflate within-subject similarity across regions. We exclude SVCCA and PWCCA from the brain analyses, as their alignment values were systematically inflated by the large voxel dimensionality relative to the number of stimuli. To assess recovery of anatomical–functional structure, we quantify region discriminability using the separability measure $d^{\prime}$, contrasting within-region similarity across subjects with across-region similarity. Across single metrics, we observe patterns consistent with the model results: metrics that preserve representational geometry or unit-level tuning better differentiate cortical regions, whereas metrics with more permissive mappings show weaker discrimination. Critically, integrating information across representational dimensions using Similarity Network Fusion (SNF) yields substantially stronger recovery of anatomical–functional structure than any individual metric. As shown in Figure 3, SNF achieves a mean $d^{\prime}$ of 21.45—nearly five times higher than the best-performing single measure—and consistently outperforms all baselines across evaluation criteria.
3.4 Data-driven typologies in models and brains
Having shown that integrating complementary representational similarity dimensions yields the most informative correspondence structure, we next examine the population-level organization revealed by the SNF-fused similarity matrices. Rather than evaluating individual pairwise comparisons, this analysis asks whether integrated similarity supports the discovery of coherent structure across populations of systems, artificial models and biological brain regions alike, without relying on prior assumptions about architecture, training paradigm, or anatomy.
3.4.1 Population structure in vision models
@@
As a baseline, clustering based on individual similarity metrics produces fragmented and inconsistent organization (Figure E.1, E.2, E.3, E.4, E.6, E.7 E.9 E.10). Metrics such as Linear Predictivity and Procrustes yield relatively uniform similarity values across models, resulting in diffuse clusters with weak separation. SoftMatch, which emphasizes unit-level geometric alignment, struggles to clearly organize models that fall outside standard CNN or supervised regimes. CCA-based metrics (PWCCA, SVCCA) produce noisy similarity matrices in which no stable clustering structure is apparent. RSA reveals partial organization (e.g. a coarse separation between CNNs and Transformers) but the resulting clusters remain diffuse. Together, these results underscore the limitations of single similarity metrics: each emphasizes a particular representational facet, leading to population structures that are incomplete or unstable.
@@
In contrast, hierarchical clustering of the fused similarity matrix reveals well-defined groupings that both recover expected relationships and reveal additional organizational structure (Figure 4, E.5, E.8, E.11)). The fidelity of these clusterings to the underlying similarity structure is quantified using the cophenetic correlation coefficient (Figure F.1), which shows that the SNF-based organization more faithfully reflects pairwise similarities than individual metrics.
Figure 4: SNF-based clustering reveals that models naturally group by architecture and supervision regime. The heatmap shows the SNF-fused similarity matrix reordered by leaf ordering. Leaf labels are colored by the cluster (formed by SNF) they belong to.
@@
Examining these clusters reveals a nuanced interplay between architecture and training paradigm. While some groupings align with architectural expectations, such as supervised CNNs and supervised ViTs forming distinct clusters (Figure 4), the most striking result is that training paradigm can override architectural boundaries. All self-supervised models, regardless of architecture, form a unified cluster that transcends the CNN–Transformer divide. Self-supervised ResNets group more closely with self-supervised ViTs than with their supervised architectural counterparts, suggesting that self-supervised objectives may induce shared computational strategies that dominate over architectural differences. Perhaps most surprisingly, hybrid architectures (ConvNeXt and Swin) cluster with MAE models despite their different design philosophies. ConvNeXts modernize CNNs with Transformer-inspired components, Swins introduce CNN-like biases into Transformers, and MAE employs masked reconstruction - yet they all converge on similar representational structures. This convergence suggests that architectural modernization and masked reconstruction, though distinct in implementation, arrive at related computational solutions.
The typology reveals the “species” of vision models, defined not just by their surface characteristics but by how they fundamentally process and organize information.
3.4.2 Population structure in visual cortex
Figure 5: PCA visualization and hierarchial score comparison of cross-brain-region representational relationships in the ventral visual pathway of the NSD Dataset. Each panel corresponds to one metric, points represent brain region instances, and text labels indicate centroid positions. The top-right bar chart displays the raw hierarchical scores for each metric, with colors corresponding to metrics.
We observe a closely parallel pattern in neural data. Applying the same SNF-based clustering approach to representational similarities across visual cortical regions reveals coherent, cross-subject organization consistent with known anatomical and functional structure. Visualization using PCA shows that while individual metrics recover broad trends, the SNF-fused similarity matrix produces the clearest and most interpretable layout, with early visual areas (V1–V4) forming a smooth progression and V4 occupying a transitional position between early and higher-level visual regions (Figure 5).
To quantify how well a representational similarity measure preserves the expected ventral-stream hierarchy, following Thobani et al. (2025), we compute a hierarchy alignment score. This score measures the rank correlation (Kendall’s $\tau$) between representational similarity and approximated hierarchical distance between cortical regions. Each region $r$ is assigned a discrete hierarchical level $L(r)\in\{1,\dots,5\}$, as $L(\text{V1v})=L(\text{V1d})=1,L(\text{V2v})=L(\text{V2d})=2,L(\text{V3v})=L(\text{V3d})=3,L(\text{V4})=4,L(\text{ventral\_visual})=5$. Given a brain-region-by-brain-region similarity matrix $S\in\mathbb{R}^{R\times R}$ and the corresponding region labels $\{r_{1},\dots,r_{R}\}$, we first construct a hierarchy-distance matrix, $H_{ij}=\bigl|L(r_{i})-L(r_{j})\bigr|,\quad 1\leq i,j\leq R.$ We then extract all off-diagonal entries of $H$ and $S$, $\mathbf{h}=\{H_{ij}:i\neq j\},\quad\mathbf{s}=\{S_{ij}:i\neq j\},$ and define $\mathbf{y}=-\mathbf{s}$ so that smaller hierarchy distance corresponds to larger similarity. The hierarchy score is the Kendall’s $\tau$-b rank correlation between $\mathbf{h}$ and $\mathbf{y}$: $\tau_{\text{hier}}=\tau_{\text{Kendall}}\!\bigl(\mathbf{h},\mathbf{y}\bigr)$. The results further support the superior performance of SNF in revealing known biological correspondence structure (Figure 5). The ability of individual similarity metrics to recover this hierarchical alignment mirrors the trends observed in the separability analyses: metrics that preserve representational geometry (e.g., RSA, linear CKA) achieve the strongest alignment with the ventral-stream hierarchy, while mapping-based metrics exhibit progressively weaker alignment as the allowable transformations become more flexible, with SoftMatch (which imposes the strictest correspondence) yielding the highest hierarchy alignment among mapping-based approaches.